OpenStudy (anonymous):

The perimeter of a rectangle is 90 ft. The ratio of its length to its width is 8 : 7. What are the dimensions of the rectangle?

OpenStudy (anonymous):

Which expresses 2700 : 4000 as a decimal?

OpenStudy (lieutenantgeneral):

Length = 8x Width = 7x Perimeter = 2*length + 2*width 90ft = 2*8x + 2*7x 90 ft = 16x + 14x 90ft = 30x 30x = 90ft x = 3ft Length = 3*8 = 24ft Width = 3*7 = 21ft. Answer: 24ft by 21 ft